a boy throws a ball upwards with an angle of 45 degrees of velocity 100 m/s that makes a semicircle in the sky. what will be the max. height of the ball????

Grade:9

5 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Shivam

initial upward velocity u=100/√2 m/s

at highest point velocity is zero

using v2=u2-2gH

put v=0

H=u2/2g = 10000/(2x2x10)= 250m

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

vikas askiitian expert
13 years ago

range of projectile = u2sin2@/g = u2/g             (at @=45)

it forms a semicircle in sky so this range is equal to diameter of this circle...

now , D = u2/g

D =2R                                          (R is radius of semicircle)

R = u2/2g

maximum height will be radius of this semicircle

Hmax = u2/2g=500m

Anirud Thyagharajan
13 years ago

hello ,

since the formula for height of a projectile is u2sin2angle/2g, the answer can be calculated thus...

bhargav teja
13 years ago

since it makes a semi circle in skymax height is half of i t's range 

  range R=U2SIN2(ANGLE)/g

     100*100/10

=1000

         then its max.height =500

Fawz Naim
13 years ago

the ball will undergo a projectile motion. the formula for the max. height attained by the ball is =u^2cos^theta/2g

theta=45 deg

u=100m/s  g=10m/s^2

therefore

H=10000*0.5/2*10

=250m